OpenStudy (ujjwal):

In what time the direction of final velocity of a projectile projected with velocity u at an angle \(\alpha\) be perpendicular to the direction of initial velocity? Ans:\[\frac{u}{g \sin \alpha}\]

OpenStudy (anonymous):

form the equation of projectile's motion. find X(t) & Y(t) and then find Y(x). it will be an equation of parabola. differentiate the equation wrt X to find dy/dx. if a is angle of projection then put dy/dx = -1/ tan(a). solve for t thereafter.

OpenStudy (ujjwal):

Thanks @quarkine

OpenStudy (foolaroundmath):

The initial velocity vector is u = ucos(a) i + usin(a) j Let the velocity at which it becomes perpendicular to u be v. Now, v = ucos(a) i + x j (x-component of velocity remains constant in projectile motion) Two vectors a and b are perpendicular if a.b = 0 (dot product) Take dot product of u and v and equate with 0 \[u^{2}\cos^{2}\alpha + u.\sin\alpha.x = 0\] \[x = -ucos^{2}\alpha/\sin\alpha\] Now, x = usin(a) - gt => t = (usin(a) - x)/g \[t = \frac{usin\alpha + \frac{ucos^{2}\alpha}{\sin\alpha}}{g}\] \[t = \frac{u}{gsin\alpha}\]

OpenStudy (ujjwal):

Thanks @FoolAroundMath ... nicely done buddy!